\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{Chetyrkin:1981qh}
K.G.~Chetyrkin and F.V.~Tkachov, \emph{{Integration by Parts: The Algorithm to Calculate beta Functions in 4 Loops}}, \href{https://doi.org/10.1016/0550-3213(81)90199-1}{\emph{Nucl. Phys. B} {\bfseries 192} (1981) 159}.

\bibitem{Kotikov:1990kg}
A.V.~Kotikov, \emph{{Differential equations method: New technique for massive Feynman diagrams calculation}}, \href{https://doi.org/10.1016/0370-2693(91)90413-K}{\emph{Phys. Lett. B} {\bfseries 254} (1991) 158}.

\bibitem{Kotikov:1991pm}
A.V.~Kotikov, \emph{{Differential equation method: The Calculation of N point Feynman diagrams}}, \href{https://doi.org/10.1016/0370-2693(91)90536-Y}{\emph{Phys. Lett. B} {\bfseries 267} (1991) 123}.

\bibitem{Gehrmann:1999as}
T.~Gehrmann and E.~Remiddi, \emph{{Differential equations for two loop four point functions}}, \href{https://doi.org/10.1016/S0550-3213(00)00223-6}{\emph{Nucl. Phys. B} {\bfseries 580} (2000) 485} [\href{https://arxiv.org/abs/hep-ph/9912329}{{\ttfamily hep-ph/9912329}}].

\bibitem{Bern:1993kr}
Z.~Bern, L.J.~Dixon and D.A.~Kosower, \emph{{Dimensionally regulated pentagon integrals}}, \href{https://doi.org/10.1016/0550-3213(94)90398-0}{\emph{Nucl. Phys. B} {\bfseries 412} (1994) 751} [\href{https://arxiv.org/abs/hep-ph/9306240}{{\ttfamily hep-ph/9306240}}].

\bibitem{Henn:2013pwa}
J.M.~Henn, \emph{{Multiloop integrals in dimensional regularization made simple}}, \href{https://doi.org/10.1103/PhysRevLett.110.251601}{\emph{Phys. Rev. Lett.} {\bfseries 110} (2013) 251601} [\href{https://arxiv.org/abs/1304.1806}{{\ttfamily 1304.1806}}].

\bibitem{Chen:1977oja}
K.-T.~Chen, \emph{{Iterated path integrals}}, \href{https://doi.org/10.1090/S0002-9904-1977-14320-6}{\emph{Bull. Am. Math. Soc.} {\bfseries 83} (1977) 831}.

\bibitem{Goncharov:1998kja}
A.B.~Goncharov, \emph{{Multiple polylogarithms, cyclotomy and modular complexes}}, \href{https://doi.org/10.4310/MRL.1998.v5.n4.a7}{\emph{Math. Res. Lett.} {\bfseries 5} (1998) 497} [\href{https://arxiv.org/abs/1105.2076}{{\ttfamily 1105.2076}}].

\bibitem{Gaiotto:2011dt}
D.~Gaiotto, J.~Maldacena, A.~Sever and P.~Vieira, \emph{{Pulling the straps of polygons}}, \href{https://doi.org/10.1007/JHEP12(2011)011}{\emph{JHEP} {\bfseries 12} (2011) 011} [\href{https://arxiv.org/abs/1102.0062}{{\ttfamily 1102.0062}}].

\bibitem{Dixon:2011pw}
L.J.~Dixon, J.M.~Drummond and J.M.~Henn, \emph{{Bootstrapping the three-loop hexagon}}, \href{https://doi.org/10.1007/JHEP11(2011)023}{\emph{JHEP} {\bfseries 11} (2011) 023} [\href{https://arxiv.org/abs/1108.4461}{{\ttfamily 1108.4461}}].

\bibitem{Dixon:2011nj}
L.J.~Dixon, J.M.~Drummond and J.M.~Henn, \emph{{Analytic result for the two-loop six-point NMHV amplitude in N=4 super Yang-Mills theory}}, \href{https://doi.org/10.1007/JHEP01(2012)024}{\emph{JHEP} {\bfseries 01} (2012) 024} [\href{https://arxiv.org/abs/1111.1704}{{\ttfamily 1111.1704}}].

\bibitem{Brandhuber:2012vm}
A.~Brandhuber, G.~Travaglini and G.~Yang, \emph{{Analytic two-loop form factors in N=4 SYM}}, \href{https://doi.org/10.1007/JHEP05(2012)082}{\emph{JHEP} {\bfseries 05} (2012) 082} [\href{https://arxiv.org/abs/1201.4170}{{\ttfamily 1201.4170}}].

\bibitem{Caron-Huot:2019vjl}
S.~Caron-Huot, L.J.~Dixon, F.~Dulat, M.~von Hippel, A.J.~McLeod and G.~Papathanasiou, \emph{{Six-Gluon amplitudes in planar $ \mathcal{N} $ = 4 super-Yang-Mills theory at six and seven loops}}, \href{https://doi.org/10.1007/JHEP08(2019)016}{\emph{JHEP} {\bfseries 08} (2019) 016} [\href{https://arxiv.org/abs/1903.10890}{{\ttfamily 1903.10890}}].

\bibitem{Caron-Huot:2011dec}
S.~Caron-Huot and S.~He, \emph{{Jumpstarting the All-Loop S-Matrix of Planar N=4 Super Yang-Mills}}, \href{https://doi.org/10.1007/JHEP07(2012)174}{\emph{JHEP} {\bfseries 07} (2012) 174} [\href{https://arxiv.org/abs/1112.1060}{{\ttfamily 1112.1060}}].

\bibitem{Golden:2013xva}
J.~Golden, A.B.~Goncharov, M.~Spradlin, C.~Vergu and A.~Volovich, \emph{{Motivic Amplitudes and Cluster Coordinates}}, \href{https://doi.org/10.1007/JHEP01(2014)091}{\emph{JHEP} {\bfseries 01} (2014) 091} [\href{https://arxiv.org/abs/1305.1617}{{\ttfamily 1305.1617}}].

\bibitem{Panzer:2014caa}
E.~Panzer, \emph{{Algorithms for the symbolic integration of hyperlogarithms with applications to Feynman integrals}}, \href{https://doi.org/10.1016/j.cpc.2014.10.019}{\emph{Comput. Phys. Commun.} {\bfseries 188} (2015) 148} [\href{https://arxiv.org/abs/1403.3385}{{\ttfamily 1403.3385}}].

\bibitem{Dennen:2015bet}
T.~Dennen, M.~Spradlin and A.~Volovich, \emph{{Landau Singularities and Symbology: One- and Two-loop MHV Amplitudes in SYM Theory}}, \href{https://doi.org/10.1007/JHEP03(2016)069}{\emph{JHEP} {\bfseries 03} (2016) 069} [\href{https://arxiv.org/abs/1512.07909}{{\ttfamily 1512.07909}}].

\bibitem{Caron-Huot:2016owq}
S.~Caron-Huot, L.J.~Dixon, A.~McLeod and M.~von Hippel, \emph{{Bootstrapping a Five-Loop Amplitude Using Steinmann Relations}}, \href{https://doi.org/10.1103/PhysRevLett.117.241601}{\emph{Phys. Rev. Lett.} {\bfseries 117} (2016) 241601} [\href{https://arxiv.org/abs/1609.00669}{{\ttfamily 1609.00669}}].

\bibitem{Mago:2020kmp}
J.~Mago, A.~Schreiber, M.~Spradlin and A.~Volovich, \emph{{Symbol alphabets from plabic graphs}}, \href{https://doi.org/10.1007/JHEP10(2020)128}{\emph{JHEP} {\bfseries 10} (2020) 128} [\href{https://arxiv.org/abs/2007.00646}{{\ttfamily 2007.00646}}].

\bibitem{Abreu:2021vhb}
S.~Abreu, R.~Britto, C.~Duhr, E.~Gardi and J.~Matthew, \emph{{The diagrammatic coaction beyond one loop}}, \href{https://doi.org/10.1007/JHEP10(2021)131}{\emph{JHEP} {\bfseries 10} (2021) 131} [\href{https://arxiv.org/abs/2106.01280}{{\ttfamily 2106.01280}}].

\bibitem{Gong:2022erh}
J.~Gong and E.Y.~Yuan, \emph{{Towards analytic structure of Feynman parameter integrals with rational curves}}, \href{https://doi.org/10.1007/JHEP10(2022)145}{\emph{JHEP} {\bfseries 10} (2022) 145} [\href{https://arxiv.org/abs/2206.06507}{{\ttfamily 2206.06507}}].

\bibitem{Yang:2022gko}
Q.~Yang, \emph{{Schubert problems, positivity and symbol letters}}, \href{https://doi.org/10.1007/JHEP08(2022)168}{\emph{JHEP} {\bfseries 08} (2022) 168} [\href{https://arxiv.org/abs/2203.16112}{{\ttfamily 2203.16112}}].

\bibitem{He:2021non}
S.~He, Z.~Li and Q.~Yang, \emph{{Truncated cluster algebras and Feynman integrals with algebraic letters}}, \href{https://doi.org/10.1007/JHEP12(2021)110}{\emph{JHEP} {\bfseries 12} (2021) 110} [\href{https://arxiv.org/abs/2106.09314}{{\ttfamily 2106.09314}}].

\bibitem{He:2021eec}
S.~He, Z.~Li and Q.~Yang, \emph{{Kinematics, cluster algebras and Feynman integrals}},  \href{https://arxiv.org/abs/2112.11842}{{\ttfamily 2112.11842}}.

\bibitem{He:2022tph}
S.~He, J.~Liu, Y.~Tang and Q.~Yang, \emph{{The symbology of Feynman integrals from twistor geometries}},  \href{https://arxiv.org/abs/2207.13482}{{\ttfamily 2207.13482}}.

\bibitem{He:2023qld}
S.~He and Y.~Tang, \emph{{Jumpstarting (elliptic) symbol integrations for loop integrals}},  \href{https://arxiv.org/abs/2304.01776}{{\ttfamily 2304.01776}}.

\bibitem{Arkani-Hamed:2017ahv}
N.~Arkani-Hamed and E.Y.~Yuan, \emph{{One-Loop Integrals from Spherical Projections of Planes and Quadrics}},  \href{https://arxiv.org/abs/1712.09991}{{\ttfamily 1712.09991}}.

\bibitem{Abreu:2017enx}
S.~Abreu, R.~Britto, C.~Duhr and E.~Gardi, \emph{{Algebraic Structure of Cut Feynman Integrals and the Diagrammatic Coaction}}, \href{https://doi.org/10.1103/PhysRevLett.119.051601}{\emph{Phys. Rev. Lett.} {\bfseries 119} (2017) 051601} [\href{https://arxiv.org/abs/1703.05064}{{\ttfamily 1703.05064}}].

\bibitem{Abreu:2017mtm}
S.~Abreu, R.~Britto, C.~Duhr and E.~Gardi, \emph{{Diagrammatic Hopf algebra of cut Feynman integrals: the one-loop case}}, \href{https://doi.org/10.1007/JHEP12(2017)090}{\emph{JHEP} {\bfseries 12} (2017) 090} [\href{https://arxiv.org/abs/1704.07931}{{\ttfamily 1704.07931}}].

\bibitem{Chen:2022fyw}
J.~Chen, C.~Ma and L.L.~Yang, \emph{{Alphabet of one-loop Feynman integrals *}}, \href{https://doi.org/10.1088/1674-1137/ac6e37}{\emph{Chin. Phys. C} {\bfseries 46} (2022) 093104} [\href{https://arxiv.org/abs/2201.12998}{{\ttfamily 2201.12998}}].

\bibitem{Dlapa:2023cvx}
C.~Dlapa, M.~Helmer, G.~Papathanasiou and F.~Tellander, \emph{{Symbol Alphabets from the Landau Singular Locus}},  \href{https://arxiv.org/abs/2304.02629}{{\ttfamily 2304.02629}}.

\bibitem{Jiang:2024eaj}
X.~Jiang, J.~Liu, X.~Xu and L.L.~Yang, \emph{{Symbol letters of Feynman integrals from Gram determinants}},  \href{https://arxiv.org/abs/2401.07632}{{\ttfamily 2401.07632}}.

\bibitem{Chen:2020uyk}
J.~Chen, X.~Jiang, X.~Xu and L.L.~Yang, \emph{{Constructing canonical Feynman integrals with intersection theory}}, \href{https://doi.org/10.1016/j.physletb.2021.136085}{\emph{Phys. Lett. B} {\bfseries 814} (2021) 136085} [\href{https://arxiv.org/abs/2008.03045}{{\ttfamily 2008.03045}}].

\bibitem{Chen:2022lzr}
J.~Chen, X.~Jiang, C.~Ma, X.~Xu and L.L.~Yang, \emph{{Baikov representations, intersection theory, and canonical Feynman integrals}}, \href{https://doi.org/10.1007/JHEP07(2022)066}{\emph{JHEP} {\bfseries 07} (2022) 066} [\href{https://arxiv.org/abs/2202.08127}{{\ttfamily 2202.08127}}].

\bibitem{Chicherin:2018old}
D.~Chicherin, T.~Gehrmann, J.M.~Henn, P.~Wasser, Y.~Zhang and S.~Zoia, \emph{{All Master Integrals for Three-Jet Production at Next-to-Next-to-Leading Order}}, \href{https://doi.org/10.1103/PhysRevLett.123.041603}{\emph{Phys. Rev. Lett.} {\bfseries 123} (2019) 041603} [\href{https://arxiv.org/abs/1812.11160}{{\ttfamily 1812.11160}}].

\bibitem{Bern:2014kca}
Z.~Bern, E.~Herrmann, S.~Litsey, J.~Stankowicz and J.~Trnka, \emph{{Logarithmic Singularities and Maximally Supersymmetric Amplitudes}}, \href{https://doi.org/10.1007/JHEP06(2015)202}{\emph{JHEP} {\bfseries 06} (2015) 202} [\href{https://arxiv.org/abs/1412.8584}{{\ttfamily 1412.8584}}].

\bibitem{Henn:2020lye}
J.~Henn, B.~Mistlberger, V.A.~Smirnov and P.~Wasser, \emph{{Constructing d-log integrands and computing master integrals for three-loop four-particle scattering}}, \href{https://doi.org/10.1007/JHEP04(2020)167}{\emph{JHEP} {\bfseries 04} (2020) 167} [\href{https://arxiv.org/abs/2002.09492}{{\ttfamily 2002.09492}}].

\bibitem{Dlapa:2021qsl}
C.~Dlapa, X.~Li and Y.~Zhang, \emph{{Leading singularities in Baikov representation and Feynman integrals with uniform transcendental weight}}, \href{https://doi.org/10.1007/JHEP07(2021)227}{\emph{JHEP} {\bfseries 07} (2021) 227} [\href{https://arxiv.org/abs/2103.04638}{{\ttfamily 2103.04638}}].

\bibitem{Baikov:1996iu}
P.A.~Baikov, \emph{{Explicit solutions of the multiloop integral recurrence relations and its application}}, \href{https://doi.org/10.1016/S0168-9002(97)00126-5}{\emph{Nucl. Instrum. Meth. A} {\bfseries 389} (1997) 347} [\href{https://arxiv.org/abs/hep-ph/9611449}{{\ttfamily hep-ph/9611449}}].

\bibitem{Chen:2023kgw}
J.~Chen, B.~Feng and L.L.~Yang, \emph{{Intersection theory rules symbology}}, \href{https://doi.org/10.1007/s11433-023-2239-8}{\emph{Sci. China Phys. Mech. Astron.} {\bfseries 67} (2024) 221011} [\href{https://arxiv.org/abs/2305.01283}{{\ttfamily 2305.01283}}].

\bibitem{Mastrolia:2018uzb}
P.~Mastrolia and S.~Mizera, \emph{{Feynman Integrals and Intersection Theory}}, \href{https://doi.org/10.1007/JHEP02(2019)139}{\emph{JHEP} {\bfseries 02} (2019) 139} [\href{https://arxiv.org/abs/1810.03818}{{\ttfamily 1810.03818}}].

\bibitem{Frellesvig:2019uqt}
H.~Frellesvig, F.~Gasparotto, M.K.~Mandal, P.~Mastrolia, L.~Mattiazzi and S.~Mizera, \emph{{Vector Space of Feynman Integrals and Multivariate Intersection Numbers}}, \href{https://doi.org/10.1103/PhysRevLett.123.201602}{\emph{Phys. Rev. Lett.} {\bfseries 123} (2019) 201602} [\href{https://arxiv.org/abs/1907.02000}{{\ttfamily 1907.02000}}].

\bibitem{Caron-Huot:2021xqj}
S.~Caron-Huot and A.~Pokraka, \emph{{Duals of Feynman integrals. Part I. Differential equations}}, \href{https://doi.org/10.1007/JHEP12(2021)045}{\emph{JHEP} {\bfseries 12} (2021) 045} [\href{https://arxiv.org/abs/2104.06898}{{\ttfamily 2104.06898}}].

\bibitem{Caron-Huot:2021iev}
S.~Caron-Huot and A.~Pokraka, \emph{{Duals of Feynman Integrals. Part II. Generalized unitarity}}, \href{https://doi.org/10.1007/JHEP04(2022)078}{\emph{JHEP} {\bfseries 04} (2022) 078} [\href{https://arxiv.org/abs/2112.00055}{{\ttfamily 2112.00055}}].

\bibitem{Giroux:2022wav}
M.~Giroux and A.~Pokraka, \emph{{Loop-by-loop differential equations for dual (elliptic) Feynman integrals}}, \href{https://doi.org/10.1007/JHEP03(2023)155}{\emph{JHEP} {\bfseries 03} (2023) 155} [\href{https://arxiv.org/abs/2210.09898}{{\ttfamily 2210.09898}}].

\bibitem{De:2023xue}
S.~De and A.~Pokraka, \emph{{Cosmology meets cohomology}}, \href{https://doi.org/10.1007/JHEP03(2024)156}{\emph{JHEP} {\bfseries 03} (2024) 156} [\href{https://arxiv.org/abs/2308.03753}{{\ttfamily 2308.03753}}].

\bibitem{Mizera:2019ose}
S.~Mizera, \emph{{Status of Intersection Theory and Feynman Integrals}}, \href{https://doi.org/10.22323/1.383.0016}{\emph{PoS} {\bfseries MA2019} (2019) 016} [\href{https://arxiv.org/abs/2002.10476}{{\ttfamily 2002.10476}}].

\bibitem{Chestnov:2022xsy}
V.~Chestnov, H.~Frellesvig, F.~Gasparotto, M.K.~Mandal and P.~Mastrolia, \emph{{Intersection Numbers from Higher-order Partial Differential Equations}},  \href{https://arxiv.org/abs/2209.01997}{{\ttfamily 2209.01997}}.

\end{thebibliography}\endgroup

\documentclass[a4paper,11pt]{article}
\usepackage{pos}
\usepackage{amsmath,latexsym,bbm}
\usepackage{braket}
\usepackage{bm}
\usepackage{booktabs}
\usepackage{setspace}
\usepackage{hepunits}
\usepackage{graphicx}% Include figure files

\usepackage{mciteplus}

\newcommand{\nn}{\nonumber}
\newcommand{\nd}{\mathrm{d}}
\newcommand{\nD}{\mathrm{D}}
\newcommand{\nT}{\mathrm{T}}
\newcommand{\nW}{\mathrm{W}}
\newcommand{\nds}{\hat{\mathrm{d}}}

\newcommand{\ep}{\epsilon}

\allowdisplaybreaks

\DeclareMathOperator{\Res}{Res}

\title{Selection rules of canonical differential equations from Intersection theory}
\author*[a]{Jiaqi Chen}
\affiliation[a]{Beijing Computational Science Research Center,\\
  Beijing 100084, China}

\emailAdd{jiaqichen@csrc.ac.cn}
\abstract{The matrix of canonical differential equations consists of the 1-$\mathrm{d}\log$-form coefficients obtained by projecting ($n$+1)-$\mathrm{d}\log$-forms onto $n$-$\mathrm{d}\log$-form master integrands. With dual form in relative cohomology, the intersection number can be used to achieve the projection and provide the selection rules for canonical differential equations, which relate to the pole structure of the $\mathrm{d}\log$ master integrands.}

\FullConference{Loops and Legs in Quantum Field Theory (LL2024)\\
 14-19, April, 2024\\
Wittenberg, Germany\\}

\begin{document}
\maketitle

\section{Introduction}
 Integration-By-Parts (IBP) \cite{Chetyrkin:1981qh} and IBP-based differential equations \cite{Kotikov:1990kg,Kotikov:1991pm,Gehrmann:1999as,Bern:1993kr} are two of the most powerful methods in perturbative quantum field theory. By appropriately selecting the master integrals, their differential equations can be transformed into a canonical form \cite{Henn:2013pwa}:
 \begin{align}
 &\nd f_I = \big( \nd \Omega \big)_{IK} f_K \, , \quad
 \big( \nd \Omega \big)_{IK} = \ep \sum_i \text{C}^{(i)}_{IK} ~ \nd \log \text{W}^{(i)}(\bm{s})  \, .  \label{eq:cde}
 \end{align} 
 In this expression, the $\big( \nd \Omega \big)_{IK}$ is a $\nd \log$-form proportional to 
 $\ep$. This property allows for the analytical solution of master integrals as multi-polylogarithm \cite{Chen:1977oja, Goncharov:1998kja}. This method  called Canonical Differential Equations (CDE)  has become a crucial analytic tool for evaluating Feynman integrals. The $\nd \log \text{W}^{(i)}(\bm{s})$ in \eqref{eq:cde} are referred to as symbol letters, and their complete set is known as the symbol alphabet. This set constrains the function space of analytic solutions and can be utilized for bootstrap techniques \cite{Gaiotto:2011dt, Dixon:2011pw, Dixon:2011nj, Brandhuber:2012vm, Caron-Huot:2019vjl}. Therefore, the properties of symbol letters, or $\nd \log \text{W}^{(i)}(\bm{s})$ in \eqref{eq:cde}, are important and have been studied extensively in \cite{Caron-Huot:2011dec, Golden:2013xva, Panzer:2014caa, Dennen:2015bet, Caron-Huot:2016owq, Mago:2020kmp, Abreu:2021vhb, Gong:2022erh, Yang:2022gko, He:2021non, He:2021eec, He:2022tph, He:2023qld, Arkani-Hamed:2017ahv, Abreu:2017enx, Abreu:2017mtm, Chen:2022fyw, Dlapa:2023cvx,Jiang:2024eaj}.

One way to 'appropriately select the master integrals' to obtain CDE is by constructing $\nd \log$ integrand as master integrands. Such construction is also related to leading singularities \cite{Henn:2013pwa, Chen:2020uyk, Chen:2022lzr, Chicherin:2018old, Bern:2014kca, Henn:2020lye, Dlapa:2021qsl}. In \cite{Chicherin:2018old,Chen:2020uyk, Chen:2022lzr}, it has been shown that Baikov representation  \cite{Baikov:1996iu} makes the construction of master integrands more convenient. Recently, \cite{Chen:2023kgw} showed how CDE emerge from $\nd \log$-form integrand of master integrals via intersection theory \cite{Mastrolia:2018uzb, Frellesvig:2019uqt}. In that work, one  only needs to use leading-order (LO) contribution and next-to-leading-order (NLO) contribution of intersection number. These contributions have universal formulas, and based on them, selection rules for CDE have been provided. However, in \cite{Chen:2023kgw}, the analysis requires the use of regulators to compute the intersection numbers, which are supposed to be set to zero at the end. The inclusion of these regulators introduces redundancies in both the calculations and the conclusions.

In this paper, we enhance the selection rules for CDE presented in \cite{Chen:2023kgw} in two aspects. Firstly, we present a new perspective on CDE by showing that the CDE matrix is the 1-$\mathrm{d}\log$-form coefficients obtained by projecting ($n$+1)-$\mathrm{d}\log$-forms onto $n$-$\mathrm{d}\log$-forms. This projection only requires the LO contribution to the intersection number, in contrast to the LO and NLO contributions used in \cite{Chen:2023kgw}. Secondly, we use the mathematical tool of twisted relative cohomology \cite{Caron-Huot:2021xqj,Caron-Huot:2021iev,Giroux:2022wav,De:2023xue} to avoid regulators. This approach helps us to    eliminate redundancy from the start. As a result, we obtain a more precise version of the CDE selection rules beyond \cite{Chen:2023kgw}.

\section{Notations of Intersection Theory and Relative Cohomology}
Intersection theory describes integrals of the form:
$ I[u,\varphi] \equiv \int  u \, \varphi \,, %\quad I_{\varphi_R} \equiv \int  u^{-1} \, \varphi_R \,,
$
where
\begin{align}
&\varphi \equiv \hat{\varphi}(\bm{z}) \bigwedge_j^n \nd z_j =\frac{Q(\bm{z})}{\left(\prod_k D_k^{a_k} \right) \left(\prod_i P_i^{b_i} \right)} \bigwedge_j^n \nd z_j \, , \quad  u=\prod_i \left[P_i(\bm{z})\right]^{\beta_i}  \, .
\label{eq:baikov}
\end{align}
Here $P_i$, $Q$ and $D_k$ are polynomials/monomials,  $a_k, b_j\in \mathbb{N}$ and $\beta_i\in \mathbb{C}$.
$I[u,\varphi]$ encompasses various integral representations, including Baikov representation and Feynman parameterization of Feynman integrals.

Then, IBP relations can be rewrite as 
\begin{equation}
\int\sum_i \partial_{z_i}(u \xi_i) \nd z_i = \int u (\varphi+\sum_i \nabla_i \xi_i)=0 \,, \, \quad \nabla_{i} = dz_i \wedge (\partial_{z_i} + \hat{\omega}_i) \, , \quad \hat{\omega}_i \equiv \partial_{z_i} \log(u) \,,
\end{equation}
where $u\xi_i$s are arbitrary integrands  belong to the integral family but in $(n-1)$-form.
Follow this expression, one can define the IBP-equivalence classes of cocycles \cite{Mastrolia:2018uzb, Frellesvig:2019uqt, Mizera:2019ose, Chestnov:2022xsy}:
\begin{equation}
\bra{\varphi_L} \equiv \varphi_L \sim \varphi_L + \sum_i \nabla_i \xi_i  \, .  \label{eq:braket}
\end{equation}
This leads to a cohomology group. Each integral $f_0=\int u \varphi_0$  corresponds to a vector in the cohomology group’s linear space \cite{Mastrolia:2018uzb, Frellesvig:2019uqt}. The master integrals $f_i=\int u \varphi_i$ form the basis of this space:
$
\bra{\varphi_0} = \sum_i c_i \bra{\varphi_i} .
$
The coefficients $c_i$ are determined by projecting rather than using IBP:
\begin{align}
&c_I=\sum_J \braket{\varphi_0|\varphi_J'}\left(\eta^{-1}\right)_{JI} \,, \quad \eta_{IJ}\equiv \braket{\varphi_I|\varphi_J'} \, , \label{eq:proj} \\
&\braket{\varphi_L|\varphi_R} = \sum_{\bm{p}} \Res_{\bm{z} = \bm{p}} \left(\psi_L \varphi_R\right) \, , \quad \nabla_1 \cdots \nabla_n \psi_L = \varphi_L \,. \label{eq:intnum}
\end{align}
Here the $\braket{\varphi_L|\varphi_R}$ denotes the intersection number, an IBP-invariant inner product.
In \eqref{eq:intnum},  $\bm{p}$ represents the isolated points where 
$n$ hypersurfaces in $\mathcal{B}=\{P_1=0,\infty,\cdots, D_1=0,\infty,\cdots\}$ intersect. Notice that $D_k$ in \eqref{eq:intnum} represent the denominators with integer powers in $u\varphi$.  In the context of Feynman integrals, such denominators typically arise from propagators. Direct computation of \eqref{eq:intnum} can be challenging with these denominators. One approach to manage this issue is to introduce regulators $\delta_k$ for such denominators in u: 
$
u=\prod_i \left[P_i(\bm{z})\right]^{\beta_i} \prod_k D_k^{\delta_k}
$
Afterward, the regulators are set to zero to obtain the final result. This method was used in  \cite{Chen:2023kgw}, but it led to redundancies.

In this paper, we use a technique known as relative cohomology to handle these denominators. While we do not delve into the mathematical details, the basic idea is as follows: we let the dual forms  $\varphi_R$ live in the maximal cut of the denominators and denote them as $\Delta_{\hat{R}}\varphi_R$. 
Practically, this is equivalent to applying the maximal cut of the sector of $\varphi_R$ to both sides before computing the intersection number.
\begin{align}
&\braket{\varphi_L|\Delta_{\hat{R}}\varphi_R}= \braket{\varphi_{L;\hat{R}}  |\varphi_{R;\hat{R}}} = \sum_{\bm{p}_{\hat{R}}}\Res_{\bm{p}_{\hat{R}}=0} \psi_{L;\hat{R}} \varphi_{R;\hat{R}}   \, , \quad \nabla_{1;\hat{R}} \cdots \nabla_{n;\hat{R}} \psi_{L;\hat{R}} = \varphi_{L;\hat{R}} \nn\\
& \varphi_{L;\hat{R}} = \Res_{\mathcal{B}_{\hat{R}}}  \left( \frac{u\varphi_L}{u_{\hat{R}}}\right)  \, , \quad  \varphi_{R;\hat{R}} = \Res_{\mathcal{B}_{\hat{R}}}  \left( \frac{\varphi_R u_{\hat{R}}}{u} \right)      \, ,  \nn\\
&u_{\hat{R}}=u|_{\mathcal{B}_{\hat{R}}}  \, , \quad  \mathcal{B}_{\hat{R}} = \{D_{R_1}=0,D_{R_2}=0,\cdots,D_{R_n}=0  \}  \, .  \label{eq:dualform}
\end{align}
Every symbol with the subscript $\hat{R}$ denotes the cut version of the original one. In \eqref{eq:dualform}, $D_{R_i}$  
 typically represent the propagators of the sector associated with $\varphi_R$. If $\mathcal{B}_{\hat{I}}  \subseteq \mathcal{B}_{\hat{J}} $, we say that $\mathcal{B}_{\hat{I}}$ is a subsector of $\mathcal{B}_{\hat{J}}$. 

To explicitly retain the information about multivariate poles in the calculations, we choose to analyze and compute the intersection number and the associated multivariate residue using a multivariate Laurent expansion. This expansion is applied after performing a factorization transformation corresponding to each pole \cite{Chestnov:2022xsy,Chen:2023kgw}:
\begin{align}
&\nT^{(\alpha)}: z_i \to f^{(\alpha)}_i(\bm{x}^{(\alpha)}) \,, \nn\\
&\nT^{(\alpha)}\left[u \right]\equiv   u(\nT^{(\alpha)}[\bm{z}]) = \bar{u}_\alpha(\bm{x}^{(\alpha)}) \prod_i \left[ x_i^{(\alpha)} - \rho_i^{(\alpha)} \right]^{\gamma_i^{(\alpha)}}  \, .
\label{eq:degenerate-1}
\end{align}
Only after this step one can legally apply a multivariate expansion to $\varphi$ and compute the intersection number:
\begin{align}
&\varphi= \sum_{\bm{b}} \varphi^{(\bm{b})}\,, \quad  \varphi^{(\bm{b})} = C^{(\bm{b})} \bigwedge_i  \left[ x_i^{(\alpha)} - \rho_i^{(\alpha)} \right]^{b_{i}} \nd x^{(\alpha)}_i \,,   \nn\\
&\braket{\varphi_L|\varphi_R}     = \sum_{\alpha} \Res_{\bm{\rho}^{(\alpha)}} \nT^{(\alpha)}\left[ \psi_L \varphi_R\right]  = \sum_{\alpha} \Res_{\bm{\rho}^{(\alpha)}} \left[ \left( \sum_{\bm{b}_L}  \nabla_1^{-1} \cdots \nabla_n^{-1} \varphi^{(\bm{b}_L)}_L  \right) \sum_{\bm{b}_R}\varphi^{(\bm{b}_R)}_R  \right] \, . \label{eq:intnumexpand}
\end{align}
Here, $\bm{b}=(b_1,\ldots,b_n)$ represents the powers in the Laurent expansion. When $b_{L,i}+b_{R,i} = -2$, we say it gives a LO contribution to intersection number. In such case, the contribution has a simple formula:
\begin{align}
&\Res_{\bm{\rho}^{(\alpha)}} \nT^{(\alpha)}\left[   \left( \nabla_1^{-1} \cdots \nabla_n^{-1}  \varphi^{(\bm{b}_L)}_L \right)\varphi^{(\bm{b}_R)}_R \right] =
\frac{C_L^{(\bm{b}_L)} C_R^{(\bm{b}_R)}}{\tilde{\bm{\gamma}}^{(\alpha)}} \,, \nn\\
&\tilde{\bm{\gamma}}\equiv \prod_i \tilde{\gamma}_i \,, \quad \tilde{\gamma}_i^{(\alpha)} = \gamma_i^{(\alpha)} - b_{R,i} - 1  \,, \ \ \  \bm{b}_L+\bm{b}_R=\bm{-2} \,  . \label{eq:intnumdlog}
\end{align}

\section{Selection Rules of CDE}
\subsection{$\nd \log$ projection and LO intersection number}
For clarity, we use $\nd$ to denote differentiation with respect to integration variables, and $\nds$ to denote differentiation with respect to one arbitrary parameter, such as kinetic parameters or masses. We define $\nD = \nd+\nds$.

To analyze and compute the CDE using \eqref{eq:proj} and \eqref{eq:dualform}, we start with the following equation:
\begin{align}
&\big( \nds \Omega \big)_{IK} = \braket{\dot{\varphi}_{I;\hat{J}}|\varphi_{J;\hat{J}}} \big( \eta^{-1} \big)_{JK} \, , \quad \eta_{IJ}\equiv \braket{\varphi_{I;\hat{J}}|\varphi_{J;\hat{J}}} \, .
\label{eq:intnumDE}
\end{align}
Here, $\varphi_I$ in both the original space and the dual space is the same $\nd \log$ integrand, constructed as $\varphi_I = \bigwedge_{j} \nd \log \nW^{(I)}_j(\bm{z})$.
The term $\dot{\varphi}_I$ in \eqref{eq:intnumDE} is defined as  $\nds f_I\equiv\int u \dot{\varphi}_I$. Therefore, we have:
\begin{align}
\nds f_I &= \int \nds \left( u \bigwedge_j \nd \log \nW^{(I)}_j(\bm{z}) \right) 
=\int u~\nD \log u \bigwedge_j \nD \log \nW^{(I)}_j  \, , \label{eq:Dlogvarphi}
\end{align}
 where $\dot{\varphi}_I$ is an ($n$+1)-$\nD \log$-form. To compute \eqref{eq:intnumDE},  we only need to compute LO contributions of the intersection number (after factorizing the pole), which are given by  \eqref{eq:intnumdlog}.

\subsection{Condition of $\braket{\dot{\varphi}_{I;\hat{J}}|\varphi_{J;\hat{J}}}\neq 0$}

Considering \eqref{eq:intnumdlog} and that $\varphi_I$ is a $\nd \log$-form, we can derive the condition under which $\braket{\dot{\varphi}_{I;\hat{J}}|\varphi_{J;\hat{J}}}\neq 0$:
\begin{align}
&\begin{cases}
\nd \log \text{-form} &: \  b_{I,i}\geq-1 \, , \ b_{J,i}\geq-1  \\
\dot{\varphi}_{I;\hat{J}} &:\  b_{\dot{I},j} = b_{I,j}-1  \\
\braket{\dot{\varphi}_{I;\hat{J}}|\varphi_{J;\hat{J}}}\neq 0 &: \ b_{\dot{I},i}+b_{J,i} \leq -2 
\end{cases} \nn\\
&\Rightarrow \braket{\dot{\varphi}_{I;\hat{J}}|\varphi_{J;\hat{J}}}\neq 0 :\  (-2\leq b_{I,j}+b_{J,j} \leq -1) \  \& \  (b_{I,i}=b_{J,i} = -1, i\neq j) \, ,
\end{align}
for specific one $j$. 
Here we omit the subscript "${;\hat{J}}$" of $b_{*,*}$ to simplify the expression, but it should be noted that all $b_{*,*}$ here correspond to the  powers after the maximal cut of the sector $\mathcal{B}_{\hat{J}}$.
We denote two cases for $\varphi_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ sharing poles as follows:
\begin{itemize}
  \item Sharing \(n_{\hat{J}}\)-Variable Simple Pole (\(n_{\hat{J}}\)-SP): This occurs when \(b_{I,j} + b_{J,j} = -1\) for one specific \(j\) and \(b_{I,i} = b_{J,i} = -1\) for all \(i \neq j\) .
  \item Sharing \((n_{\hat{J}} - 1)\)-Variable Simple Pole (\((n_{\hat{J}} - 1)\)-SP): This occurs when \(b_{I,i} = b_{J,i} = -1\).
\end{itemize}

For the case in which they share an ($n_{\hat{J}}$-1)-SP,  $\nD \log \left( \nT^{(\alpha)} [u] \right)$ in \eqref{eq:Dlogvarphi} provides a pole of one remaining variable. The shared pole contributes
\begin{equation}
-\frac{\gamma^{(\alpha)}_{j;\hat{J}} }{\bm{\gamma}^{(\alpha)}_{\hat{J}} } \, \nds \int C_{I}^{(\bm{b}_{I} )}C_J^{(\bm{b}_J)} \, \nds \rho^{(\alpha)}_{j;{\hat{J}}} 
\label{eq:NMCSPdlog}
\end{equation}
to intersection number $\braket{\dot{\varphi}_{I;\hat{J}}|\varphi_{J;\hat{J}}}$.
It is a $\nds \log$-form as discussed in \cite{Chen:2023kgw}.

For the case in which they share an $n_{\hat{J}}$-SP, the shared pole contributes
\begin{equation}
\frac{C_I^{(\bm{-1})} C_J^{(\bm{-1})}}{\bm{\gamma}^{(\alpha)}_{\hat{J}}   } \, \nds \log \Big( \bar{u}_{\alpha;\hat{J}}(\bm{\rho}^{(\alpha)}_{\hat{J}} ) \Big)   \label{eq:CSPdlog}
\end{equation}
to intersection number $\braket{\dot{\varphi}_{I;\hat{J}}|\varphi_{J;\hat{J}}}$.

\subsection{Condition of $\big( \eta^{-1} \big)_{JK}\neq 0$}

Notice that $\eta_{JK}$ is non-zero only when $\varphi_{J;\hat{K}}$ and $\varphi_{K;\hat{K}}$ share at least an $n_{\hat{J}}$-SP and each shared $n_{\hat{J}}$-SP contributes $
C_J^{(\bm{-1})} C_K^{(\bm{-1})}/\bm{\gamma}^{(\alpha)}_{\hat{K}}  
$. Also, note that $\big( \eta^{-1} \big)_{JK} = (-1)^{J+K} |\eta^{(KJ)}| / |\eta|$, where $\eta^{(KJ)}$ is the adjoint matrix of $\eta$. Following a similar analysis as in \cite{Chen:2023kgw}, we need to modify the concept of "n-SP chain" from \cite{Chen:2023kgw}, which is unoriented, to a new concept of {\bf cut-$n$-SP} chain:
\begin{itemize}
\item If ${\varphi}_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ share $n_{\hat{J}}$-SP,  we say $\varphi_I$ is cut-$n$-SP related to $\varphi_J$, and denoted as $\varphi_I \rightarrow \varphi_J$ or $\varphi_J \leftarrow \varphi_I$. Now, it is an orient relation, i.e., $\varphi_I \rightarrow \varphi_J$ does not imply $\varphi_J \rightarrow \varphi_I$. 

\item If $\varphi_I \rightarrow \varphi_J$, we also say that $\varphi_I$ is linked to $\varphi_J$ via a cut-$n$-SP chain. If $\varphi_I \rightarrow \varphi_J$ and $\varphi_J \rightarrow \varphi_K$, then we say $\varphi_I$ is linked to $\varphi_K$ via the cut-$n$-SP chain $\varphi_I \rightarrow \varphi_J \rightarrow \varphi_K$, we denote it as $\varphi_I \rightarrow\rightarrow \varphi_J$. Similar understanding for more forms $\varphi$.  
\end{itemize}
Then, \(\big( \eta^{-1} \big)_{JK}\) could be non-zero only when \(\varphi_J \rightarrow\rightarrow \varphi_K\).

\subsection{Conclusion}
With the above discussions,  we have CDE selection rules: 
If all master integrals have $\nd \log$-form $\varphi_I$s, the differential equations $\big( \nds \Omega \big)_{IK}$  is $\nd \log$-form and given by the simple formula of  LO contribution of intersection number. When all $\beta_i$ in eq.\eqref{eq:baikov} are proportional to $\ep$, $\big( \nds \Omega \big)_{IK}$ is reduced to be proportional to $\ep$, thus is canonical form. $\big( \nds \Omega \big)_{IK}$ could be non-zero only when there exist a $\varphi_J$ that  $\dot{\varphi}_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ share $n_{\hat{J}}$-SP or ($n_{\hat{J}}-1$)-SP, and $\varphi_J \rightarrow \rightarrow \varphi_K$.

\section{Univariate Example: Comparing Regulator Method with Relative Cohomology}
Consider $u=z^\delta (z-c_1)^\ep(z-c_2)^\ep$, where  $\delta$ 
 is regulator. We construct the $\nd \log$ basis as follow:
\begin{align}
\varphi_1 = \frac{\nd z}{z} =\nd \log z \, ,  \quad \varphi_2 = \frac{\nd z}{z-c_1} - \frac{\nd z}{z-c_2}  = \nd \log \left( \frac{z-c_1}{z-c_2}\right)  \, .
\end{align}

\subsection{Regulator method}
If we compute the differential equations of the basis using the regulator method, we obtain:
\begin{align}
&\omega=\nd \log u = \left( \frac{\delta}{z} + \frac{\ep}{z-c_1}+ \frac{\ep}{z-c_2} \right) \nd z \, . \nn\\
&\mathcal{B}=\{z=0,z-c_1=0,z-c_2=0,z=\infty\}\,,\quad \{\bm{p}\}=\{0,c_1,c_2,\infty\} \, . \nn\\
&\gamma_1=\delta\, , \ \ \gamma_2=\ep\, , \ \ \gamma_3=\ep\, , \ \ \gamma_4=-2\ep-\delta \, .
\end{align}
Then, $\eta_{11}$ receives contributions from shared 1-SP at $p_1=0$ and $p_4=\infty$ of $\varphi_1$ itself, so it equals $1/\gamma_1+1/\gamma_4$. A similar analysis can be applied to other \(\eta_{JK}\), giving:
\begin{align}
&\eta=\braket{\varphi_I|\varphi_J}=\left(
\begin{array}{cc}
\frac{1}{\gamma_1}+\frac{1}{\gamma_4} & 0 \\
0 &  \frac{1}{\gamma_2}+\frac{1}{\gamma_3}  \\
\end{array}
\right)=\left(
\begin{array}{cc}
\frac{1}{\delta }-\frac{1}{\delta +2 \epsilon } & 0 \\
0 & \frac{2}{\epsilon } \\
\end{array}
\right)  \, , \nn\\
&\eta^{-1}=\left(
\begin{array}{cc}
\delta(\delta+2\ep)/(2 \epsilon)  & 0 \\
0 & \epsilon/2 \\
\end{array}
\right)  \, .
\end{align}
For $\braket{\dot{\varphi}_I|\varphi_J}$, we have $\dot{\varphi}_I$
\begin{align}
&\dot{\varphi}_1 = \nD \left(\log \bar{u}_1(z) + \delta \log z \right) \wedge \nD \log z  = \nD \log \bar{u}_1(z) \wedge \nD \log z  \, , \nn\\
&\dot{\varphi}_2 =  \nD \log \bar{u}_2(z) \wedge \nD \log (z-c_1) - \nD \log \bar{u}_3(z) \wedge \nD \log (z-c_2)   \nn\\
&\quad = \delta ~ \nD \log z \wedge \varphi_2 + 2 \ep ~ \nD \log (z-c_2) \wedge \nD \log (z-c_1)  \, , \nn\\
&\bar{u}_1(z) = (z-c_1)^\ep(z-c_2)^\ep \, , \quad  \bar{u}_2(z) = z^\delta(z-c_2)^\ep \, , \quad  \bar{u}_3(z) = z^\delta(z-c_1)^\ep \, .
\end{align}
Let's calculate $\braket{\dot{\varphi}_1|\varphi_1}$ as an example. For the contribution from the shared 1-SP $p_1=0$, we use \eqref{eq:CSPdlog} to get
\begin{align}
\frac{1}{\delta} \nds \log \left((z-c_1)^\ep(z-c_2)^\ep|_{z=0} \right) = \frac{\ep}{\delta} \nds \log(c_1c_2)  \,,
\end{align}
For the shared 1-SP at $p_4=\infty$:
\begin{align}
\frac{1}{-\delta-2\ep} \nds \log \left((1-c_1 t)^\ep(1-c_2 t)^\ep|_{t=0} \right) = \frac{\ep}{-\delta-2\ep} \nds \log 1 = 0 \,,
\end{align}
where $t=1/z$. Thus, this term does not contribute. Continuing the calculation, we find:
\begin{align}
&\Big( \braket{\dot{\varphi}_I|\varphi_J} \Big) = \nds \left(
\begin{array}{cc}
\frac{\epsilon }{\delta } \log \left(c_1c_2\right) & \log
\left(c_2/c_1\right) \\
\log \left(c_2/c_1\right) & \frac{\delta }{\epsilon } \log \left(c_1c_2\right) + 4\log(c_1-c_2) \\
\end{array}
\right)  \, .  \nn\\
&\nds \Omega^{\text{reg}} (c_i;\ep,\delta) = \nds \left(
\begin{array}{cc}
\frac{\delta +2 \ep}{2} \log \left(c_1 c_2\right)& \frac{\epsilon}{2}   \log \left(\frac{c_2}{c_1}\right) \\
\frac{\delta   (\delta +2 \epsilon )}{2 \epsilon } \log \left(\frac{c_2}{c_1}\right) 
& \frac{\delta}{2}   \log \left(c_1 c_2\right)+2 \epsilon  \log \left(c_1-c_2\right) \\
\end{array}
\right)   \, . \label{eq:1varDEreg}
\end{align}
It is redundant and $\nds \Omega^{\text{reg}}_{21}$ vanishes when $\delta=0$, but the selection rules get from the regulator method in \cite{Chen:2023kgw} do not  indicate this.

\subsection{Relative cohomology method}
If we calculate $\Omega$ using relative cohomology method, we have 
\begin{align}
&u= (z-c_1)^\ep(z-c_2)^\ep
\, , \ \  D_1=z   \nn\\
&\omega=\nd \log u = \left( \frac{\ep}{z-c_1}+ \frac{\ep}{z-c_2} \right) \nd z \, . \quad
\{\bm{p}\}=\{0,c_1,c_2,\infty\} \, .
\end{align}
Calculation of $\braket{\varphi_I|\Delta_{\hat{1}}\varphi_1}$ is simplified as we apply the cut of $\mathcal{B}_{\hat{1}}$:
\begin{align}
&\braket{\varphi_1|\Delta_{\hat{1}}\varphi_1}=\braket{\varphi_{1;\hat{1}}|\varphi_{1;\hat{1}}} = \braket{1|1}=1  \nn\\
&\braket{\varphi_2|\Delta_{\hat{1}}\varphi_1}=\braket{\varphi_{2;\hat{1}}|\varphi_{1;\hat{1}}} = \braket{0|1}=0 \nn\\
&\braket{\dot{\varphi}_1|\Delta_{\hat{1}}\varphi_1}=\braket{\dot{\varphi}_{1;\hat{1}}|\varphi_{1;\hat{1}}} = \braket{\nds\log u|_{z=0}|1}= \ep \nds \log (c_1 c_2) \nn\\
&~\dot{\varphi}_2= \nD \log u \wedge \nD \log \left(z-c_1\right) - \nD \log u \wedge \nD \log \left(z-c_2\right) \nn\\
& \quad ~ = 2 \ep ~  \nD \log \left(z-c_2\right) \wedge \nD \log \left(z-c_1\right) \, , \nn\\
&\braket{\dot{\varphi}_2|\Delta_{\hat{1}}\varphi_1}=\braket{\dot{\varphi}_{2;\hat{1}}|\varphi_{1;\hat{1}}} = \braket{0|1}= 0 \, .
\end{align}
Thus, we have
\begin{align}
&\eta=\braket{ \varphi_{I;\hat{J}}|\varphi_{J;\hat{J}}} =\left(
\begin{array}{cc}
1 & 0 \\
0 & 2/\epsilon  \\
\end{array}
\right)  \, , \quad  \eta^{-1}=\left(
\begin{array}{cc}
1  & 0 \\
0 & \epsilon/2 \\
\end{array}
\right)  \, , \nn\\
&\Big( \braket{\dot{\varphi}_{I;\hat{J}}|\varphi_{J;\hat{J}}} \Big) = \nds \left(
\begin{array}{cc}
\ep \log \left(c_1 c_2\right) 
& \ep \log \left(c_2/c_1\right) \\
0 
& 4 \log
\left(c_1-c_2\right) \\
\end{array}
\right)  \, \nn\\
&\nds \Omega =\Big( \braket{\dot{\varphi}_{I;\hat{J}}|\varphi_{J;\hat{J}}} \Big).\eta^{-1}= \ep~  \nds \left(
\begin{array}{cc}
\log \left(c_1 c_2\right) & \frac{1}{2} \log \left(c_2/c_1\right) \\
0 & 2  \log
\left(c_1-c_2\right) \\
\end{array}
\right)  \, .
\end{align}
We have $\nds \Omega=\nds \Omega^{\text{reg}} (c_i;\ep,0)$. This demonstrates that the relative cohomology method simplifies the computation, avoids redundancy from regulators, and provides a more precise version of the selection rules for CDE. In particular, it correctly shows that $\Omega_{21}=0$ in this case. 

\section{Summary and Outlook}
We improve the selection rules for CDE by providing a perspective from the $\nd \log$ projection and avoiding redundancy through the relative cohomology method. The mathematical structure we have developed can benefit various aspects of perturbative quantum field theory, including reduction (since CDEs are themselves reduction relations), symbolic bootstrap, and the analytic evaluation of Feynman integrals.

\section*{Acknowledgements}
We thank Andrzej Pokraka for valuable discussions. This work is supported by Chinese NSF funding under Grant No.11935013, No.11947301, No.12047502 (Peng Huanwu Center), No.12247120, No.12247103, NSAF grant No.U2230402, and China Postdoctoral Science Foundation No.2022M720386.

\vspace{-1mm}
\setlength{\bibsep}{2pt}
\bibliographystyle{JHEP}
\bibliography{ref}

\end{document}